OpenStudy (anonymous):

x^3+ 2x^2 - 3x + 4 = ??

OpenStudy (mathmale):

Some thought required here. First, you could assume that this graph represents a polynomial. You may (or may not) be able to determine the exact polynomial. Four dots appear on the graph, each representing a particular point (with x- and y-value). Note how you have two roots / solutions. I'd suggest you list all of the points, estimating the one for which x is approx. 1.4 (which represents a root of the polynomial). If you make that into a dot, you'll then have FIVE dots representing 5 points on the curve. you could use the model y= ax^5 + bx^4 + cx^3 + dx^2 + ex + f which has 6 unknown coefficients: { a, b, c, d, e, f }. Please read this through and then tell me which part of my comments make sense to you and which do not.

OpenStudy (mathmale):

I'm suggesting we ignore the x-intercept that is unmarked. That leaves us with FOUR dots, right? Four dots => four points on graph => Fourth order polynomial with 5 coefficients.

OpenStudy (mathmale):

Let the polynomial be \[f(x)=ax^4+bx^3+cx^2 +dx +e\]

OpenStudy (mathmale):

You'll see that the first point is (0,3). Take the function I just gave you; let f(0) = 3. This results in\[f(0)=3=a*0^4 +b*0^3+c*0^2+d*0 +e\]

OpenStudy (mathmale):

what is the value of e?

OpenStudy (mathmale):

And what was that other something that y ou were supposed to find?

OpenStudy (mathmale):

Hint: 3=0+0+0+0+e, so e=?

OpenStudy (mathmale):

sorry...OpenStudy was down for a few minutes. If you have four black dots on the given graph, that means we can find the values of only four (not five) coefficients, { a, b, c, d }. Thus, discard the fourth-order model, above, and adopt the third-order model \[f(x) = ax^3 +bx^2+cx+d\]If you now focus on the first point (the black dot at (0,3), and let f(0)=3, then d = 3. Then \[f(x)=ax^3+bx^2+cx+3.\] Now, look at the other three dots. Substitute the x and y-values into the above equation for each point. Each time you do this, you will have found the value of one more coefficient. You will end up knowing the coeff. { a, b, c, d }. Write out the polynomial using these known (numerical) coeff. instead of the lettters { a, b, c, d } to represent the coeff. Then you're done! Unfortunately, I have to get off the Internet now. But if you need further help, simply respond here, and I will see your message next time I'm on OpenStudy.

OpenStudy (mathmale):

I've returned. Need any further help / suggestions?